# fPLSA: Learning Semantic Structures in Document Collections Using Foundation Models

### v1

@@ 4.4 Baselines
We compare our approach with the traditional Latent Dirichlet Allocation (TradLDA) algorithm designed to discover latent topics in a collection of text spans (Blei et al., 2003).
@@
As Li et al. (2023) showed that the topic labels generated by LLMs based on the key terms learned through TradLDA are preferred more often than the original labels, we also include TradLDA+LLM as a baseline. Specifically, we first learn the topics with the key terms for each topic using TradLDA, and then use GPT-4 to generate a description for each topic based on the key terms.
@@
Recent work showed that, with appropriate prompts, LLMs are capable of directly generating topic labels given a set of text documents and condensing overarching topics (Mu et al., 2024). As a baseline, we adapt the approach (along with the prompts) to generate topic descriptions for each text segment.
@@ 4.5 Large Language Model Setup
For clustering and tagging, we use GPT-4 for all approaches, a powerful LLM (OpenAI et al., 2024). We set $top\_p=0.5$, sampling temperature $\tau=1.0$, zero frequency and presence penalty. We also use GPT-4 with $top\_p=0.5$ to estimate the reconstruction log-likelihood. We set the temperature $\tau=1.0$ when sampling alternative segments and $\tau=0$ when choosing the best continuation.
To measure Hits@K Accuracy, we use ChatGPT (gpt-3.5-turbo; OpenAI (2023)) instead of GPT-4, because GPT-4 may have data contamination issues (Deng et al., 2024) on MATH and BBH benchmarks based on its timestamp. We set $top\_p=0.5$ and temperature $\tau=1.0$ when sampling solutions from ChatGPT.
@@ 5.1 Reconstruction Likelihood
No Tag TradLDA TradLDA+LLM Prompting fPLSA
@@
MATH-All -3.67 -3.16 -3.57 -3.44 -3.04
Table 1: Reconstruction log-likelihood of fPLSA versus the baseline without tags (No Tag), traditional LDA (TradLDA), traditional LDA with LLM-generated tag descriptions (TradLDA+LLM(Li et al., 2023), and the prompting baseline (Prompting(Mu et al., 2024) on WritingPrompts story dataset, Number Theory dataset from MATH (MATH-Num), and the whole MATH (MATH-All) dataset.
Keywords Tag Description
@@
Table 2: Examples of keywords learned on short story segments in WritingPrompts through TradLDA and the corresponding tag descriptions generated by GPT-4. Given only the keywords without context, the tag descriptions produced by GPT-4 are too generic to recover the original text spans.
Prompting Tags fPLSA Tags
@@
Table 3: Example tags learned on short story segments in WritingPrompts through Prompting versus fPLSA. Prompting tags are either too mixed (e.g. Tag 1 and 2) or too generic (e.g. Tag 3), while fPLSA groups segments of similar themes into the same cluster and describes each cluster with detailed explanations and example plots.
First, we compare the reconstruction log-likelihood of fPLSA with the No Tag baseline (without conditioning on any tags). As shown in Table 1, conditioning on fPLSA tags helps predict the original texts: fPLSA brings 0.6–1.4 higher log-likelihood than the No Tag baseline.
TradLDA also brings higher reconstruction log-likelihood over the No Tag baseline. However, since TradLDA only captures word or term co-occurrences, it still underperforms fPLSA consistently on all three datasets. Moreover, TradLDA+LLM fails to improve over TradLDA. As shown by the examples in Table 2, it is extremely challenging for LLMs and even humans to extract meaningful semantic information from the keywords learned on short text segments through TradLDA, and the resulting tag descriptions are overly generic, making it challenging to reconstruct the original text segments accurately.
Compared with the Prompting baseline, fPLSA achieves 0.2–0.4 higher log-likelihood on all three datasets. We further compared the tags learned using Prompting versus fPLSA. As shown by the examples in Table 3, Prompting tends to merge unrelated topics into a mixed topic (e.g. Tag 1 and 2), and the resulting topics become overly broad. Even for tags sharing a common theme, the descriptions often lack specificity and detail (e.g. Tag 3). By contrast, fPLSA identifies segments with similar themes, groups them into a single cluster and produces more detailed tag descriptions with example plots.
@@ 5.2 Hits@K Accuracy
No Tag GenOutline TradLDA TradLDA+LLM Prompting fPLSA
@@
InterAlgebra 55.7 51.7 58.7 59.2 61.2 64.7
Number 65.4 76.0 77.9 74.0 78.8 78.8
PreAlgebra 74.2 79.1 81.3 81.3 84.6 83.0
PreCalculus 42.2 46.8 51.4 46.8 49.5 54.1
Average 62.9 65.6 70.1 67.6 71.1 72.3
@@
Table 4: Hits@K accuracy of fPLSA versus directly sampling without tags (No Tag), two-step sampling with LLM-generated outline (GenOutline), traditional LDA (TradLDA), traditional LDA with LLM-generated tag descriptions (TradLDA+LLM(Li et al., 2023), and the prompting baseline (Prompting(Mu et al., 2024) on 12 challenging tasks from BBH benchmark (Suzgun et al., 2022) and 7 tasks from MATH (Hendrycks et al., 2021).
TradLDA Tags
This cluster often contains words such as either, distinct, case, problem, must, find, 10, three, 72, follows, 3a, yields, since, digit, thus, digits, equal, 2a, 144, base.
This cluster often contains words such as 250, shown asy, makes, means, becomes, coordinates, sphere, origin, thus, left frac pi, frac pi right, pi right, cos frac, pi frac pi, pi pi, frac pi frac, pi frac, frac, frac pi, pi.
This cluster often contains words such as equation, note, sqrt, also, line, get, 2b, rightanglemark, abc, draw rightanglemark, 25 boxed, dfrac, must, since, let, expanding, property, 300, angle, xy.
This cluster often contains words such as 2t, makes, circ boxed, triangle, 120, 120 circ, 60 circ, 90 circ, circ angle, operatorname, 360 circ, 360, since, 45 circ, 180 circ, 90, 180, angle, 45, circ.
This cluster often contains words such as 40, also, overline, 14, therefore, bc, align therefore, end align therefore, circ, let, frac cdot, respectively, sqrt, triangle, cosines, law cosines, cdot, law, frac, angle.
Prompting Tags
Algebra and equations in manipulation and solving.
Algebraic manipulation and polynomial factorization.
Equation setup and solving for ages, distances, and quantities.
Inverse function calculations and summation.
Geometry and trigonometry in problem-solving.
fPLSA Tags
Using congruence or similarity to deduce equal angles or sides in geometric figures.
Perform algebraic manipulations to solve for an unknown variable.
Utilizes specific mathematical theorems or properties, such as De Moivre’s Theorem or the Law of Cosines, to solve problems.
Identify or prove relationships between angles, sides, or other elements of geometric figures.
This tag includes steps that conclude a mathematical procedure or finalize the simplification of an expression.
Table 5: Top 5 tags from TradLDA, Prompting and fPLSA that lead to the highest Hits@K Accuracy on MATH.
(a) BBH
(b) MATH
Figure 1: Heatmap of the average Hits@K Accuracy of tags learned mostly from a particular task when used on other tasks. The x axis represents the task from which the tags are learned from, and the y axis represents the test task. Tags learned from tasks other than the test task are proven to be helpful and sometimes even more helpful than the tags from the test task.
@@
We further evaluate how the semantic structural tags help with downstream generation by measuring the Hits@K Accuracy of various sampling methods with or without tags. First, compared with direct sampling without using any tags, hierarchical sampling with fPLSA tags leads to significantly higher Hits@K accuracy by +9.4 points on MATH and +16.6 points on BBH on average. Additionally, we compare fPLSA with GenOutline, a two-step sampling approach where we prompt the LLM to generate an outline before generating the actual solution. GenOutline improves over direct sampling on most tasks, but still underperforms hierarchical sampling with fPLSA by 7–9 points. These results indicate that hierarchical sampling using tags derived from the domain-specific documents via fPLSA produces more effective and diverse output solutions, thereby increasing the likelihood of hitting the correct answer.
@@
Next, we compare fPLSA with hierarchical sampling with existing tagging approaches. fPLSA tags lead to more diverse outputs with a higher chance of hitting the correct solution paths than TradLDA on 16 out of 19 tasks. It brings an an average accuracy improvement of 2–10 points over TradLDA. Similarly, compared with TradLDA+LLM, fPLSA achieves higher Hits@K Accuracy on 17 out of 19 tasks and improves the average accuracy by 5–11 points across BBH and MATH. Compared with the Prompting baseline, fPLSA achieves higher Hits@K Accuracy on 14 out of 19 tasks. Overall, hierarchical sampling with fPLSA tags improves Hits@K Accuracy over existing tagging approaches by 1–11 points on average.
We further examine the top 5 tags from each tagging approach that lead to the highest Hits@K Accuracy when used as part of the outline. As shown in Table 5, the TradLDA tags are too low-level, making it difficult for an LLM to follow. The Prompting tags, however, are too generic – for example, the tag “Algebra and equations in manipulation and solving” covers almost all solution steps in algebra problems. By contrast, fPLSA tags are more specific and instructive than the Prompting tags, but are still representative of groups of solution steps.
Finally, we investigate whether the tags learned through fPLSA generalize across tasks. Specifically, we examine the average Hits@K Accuracy of tags learned mostly from a particular task when used on other tasks. As shown in Figure 1, tags learned from tasks other than the test task are helpful in sampling effective solutions and sometimes even more helpful than the tags learned on the test task itself. This is possibly because the LLM is already familiar with the solution paths suggested by the tags learned from the test task itself, while the tags learned from other tasks may cause the LLM to think out of the box.
@@ 6 Conclusion
We introduced fPLSA, a foundation-model-based Probabilistic Latent Semantic Analysis method that aims to uncover the latent semantic structures in document collections by iteratively clustering and tagging document segments based on document-level contexts. Our experiments on story writing, math and multi-step reasoning tasks show that fPLSA tags are more informative in reconstructing the original texts than tags generated by existing tagging methods. fPLSA tags are also useful in generating more diverse solutions via hierarchical sampling and lead to higher Hits@K Accuracy than existing methods. These results suggest the potential of fPLSA for generating effective task guidelines given some worked-out examples, along with hierarchical sampling and searching for problem solutions based on a verification or reward model.

### v2

@@ 4.4 Baselines
We compare our approach with the traditional Latent Dirichlet Allocation (TradLDA), a type of LSA algorithm designed to discover latent topics in a collection of text spans (Blei et al., 2003).
@@
As Li et al. (2023) showed that the topic labels generated by LLMs based on the key terms learned through TradLDA are preferred more often than the original labels, we also include TradLDA+LLM as a baseline. Specifically, we first learn the topics and the key terms for each topic using TradLDA, and then use GPT-4 to generate a description for each topic based on the key terms.
@@
Recent work showed that, with appropriate prompts, LLMs are capable of directly generating topic labels given a set of text documents and condensing overarching topics (Pham et al., 2024; Wang et al., 2023; Mu et al., 2024). As a baseline, we adapt the approach (along with the prompts) in Mu et al. (2024) to generate topic descriptions for each text segment.
@@ 4.5 Large Language Model Setup
For clustering and tagging, we use GPT-4 (OpenAI et al., 2024) and Qwen-2.5-7B (a much smaller LLM introduced in Qwen et al. (2025)). We also use GPT-4 to estimate the reconstruction log-likelihood. To measure Hits@K Accuracy, we use ChatGPT (gpt-3.5-turbo; OpenAI (2023)) instead of GPT-4, because GPT-4 has achieved high accuracy on MATH and BBH (e.g. 84% on MATH (Zhou et al., 2023)), possibly due to data contamination issues (Deng et al., 2024; Bubeck et al., 2023). Thus, we use ChatGPT for solution sampling to show the potential of using learned tags to diversify the sampled outputs and improve the chance of finding a correct answer when the model cannot find it through direct sampling. [^4] More details in the Appendix.
@@ 5.1 Reconstruction Likelihood
No Tag TradLDA TradLDA+LLM Prompting fLSA
@@
MATH-All -3.67 -3.16 -3.57 -3.44 -2.94
Table 1: Reconstruction log-likelihood of fLSA versus the baseline without tags (No Tag), traditional LDA (TradLDA), traditional LDA with LLM-generated tag descriptions (TradLDA+LLM(Li et al., 2023), and the prompting baseline (Prompting(Mu et al., 2024) on WritingPrompts story dataset, Number Theory dataset from MATH (MATH-Num), and the MATH (MATH-All) dataset.
Key Terms Tag Description
@@
Table 2: Examples of key terms learned on short story segments in WritingPrompts through TradLDA and the corresponding tag descriptions generated by GPT-4. Given only the key terms without context, the tag descriptions produced by GPT-4 are too generic to recover the original text spans.
Prompting Tags fLSA Tags
@@
Table 3: Example tags learned on short story segments in WritingPrompts through Prompting versus fLSA. Prompting tags are either too mixed (e.g. Tag 1 and 2) or too generic (e.g. Tag 3), while fLSA groups segments of similar themes into the same cluster and describes each cluster with detailed explanations and example plots.
First, we compare the reconstruction log-likelihood of fLSA with the No Tag baseline (without conditioning on any tags). As shown in Table 1, conditioning on fLSA tags helps predict the original texts: fLSA brings 0.7–1.4 higher log-likelihood than the No Tag baseline.
TradLDA also brings higher reconstruction log-likelihood over the No Tag baseline. However, since TradLDA only captures word or term co-occurrences, it still underperforms fLSA consistently on all three datasets. Moreover, TradLDA+LLM fails to improve over TradLDA. As shown by the examples in Table 2, it is extremely challenging for LLMs and even humans to extract meaningful semantic information from the key terms learned on short text segments through TradLDA, and the resulting tag descriptions are overly generic, making it challenging to reconstruct the original text segments accurately.
Compared with the Prompting baseline, fLSA achieves 0.2–0.5 higher log-likelihood on all three datasets. We further compared the tags learned using Prompting versus fLSA. As shown by the examples in Table 3, Prompting tends to merge unrelated topics into a mixed topic (e.g. Tag 1 and 2), and the resulting topics become overly broad. Even for tags sharing a common theme, the descriptions often lack specificity and detail (e.g. Tag 3). By contrast, fLSA identifies segments with similar themes, groups them into a single cluster and produces more detailed tag descriptions with example plots.
@@ 5.2 Hits@K Accuracy
No Tag GenOutline TradLDA TradLDA+LLM Prompting fLSA
@@
InterAlgebra 55.7 51.7 58.7 59.2 61.2 61.2
Number 65.4 76.0 77.9 74.0 78.8 83.7
PreAlgebra 74.2 79.1 81.3 81.3 84.6 89.0
PreCalculus 42.2 46.8 51.4 46.8 49.5 55.0
Average 62.9 65.6 70.1 67.6 71.1 72.9
@@
Table 4: Hits@K accuracy of fLSA versus directly sampling without tags (No Tag), two-step sampling with LLM-generated outline (GenOutline), traditional LDA (TradLDA), traditional LDA with LLM-generated tag descriptions (TradLDA+LLM(Li et al., 2023), and the prompting baseline (Prompting(Mu et al., 2024) on 12 challenging tasks from BBH benchmark (Suzgun et al., 2022) and 7 tasks from MATH (Hendrycks et al., 2021).
@@
We further evaluate how the tags and semantic structure learned through fLSA help expand the output space in the right directions that lead to correct solutions by measuring the Hits@K Accuracy of various sampling methods with or without tags. First, compared with direct sampling without using any tags, hierarchical sampling with fLSA tags leads to significantly higher Hits@K accuracy by +10.0 points on MATH and +16.6 points on BBH on average. Additionally, we compare fLSA with GenOutline, a two-step sampling approach where we prompt the LLM to generate an outline before generating the actual solution. GenOutline improves over direct sampling on most tasks, but still underperforms hierarchical sampling with fLSA by 7–9 points. These results indicate that hierarchical sampling using tags derived from the domain-specific documents via fLSA produces more effective output solutions, thereby increasing the likelihood of hitting the correct answer with K samples.
@@
Next, we compare fLSA with hierarchical sampling with existing tagging approaches. fLSA tags expand the output space in the directions that lead to correct answers more often than TradLDA on 16 out of 19 tasks. It brings a significant improvement of 3–10 points over TradLDA. [^5] We test significance using paired student’s t-test with significance level $\alpha=0.05$. Similarly, compared with TradLDA+LLM, fLSA achieves higher Hits@K Accuracy on 17 out of 19 tasks and improves the average accuracy by 5–11 points across BBH and MATH. Compared with the Prompting baseline, fLSA achieves higher Hits@K Accuracy on 14 out of 19 tasks. Overall, hierarchical sampling with fLSA tags improves Hits@K Accuracy significantly over existing tagging approaches by 2–11 points on average.
5.3 Learning Tags with Smaller LLMs
In addition to GPT-4, we also evaluate fLSA using a smaller LLM – Qwen-2.5-7B. We run fLSA using Qwen-2.5-7B as the base model (while the other hyper-parameters remain unchanged) and measure the Hits@K Accuracy of hierarchical sampling using the learned tags on BBH. We discover that the average accuracy drops by 5 points compared to the tags learned using GPT-4, but it still outperforms TradLDA and Prompting (using the much larger GPT-4 model) by 3–6 points.
5.4 Ablation Study
We further examine how the number of tags learned through fLSA influences its ability to expand the output space. Specifically, we compare the Hits@K Accuracy of hierarchical sampling with 20, 50, and 100 fLSA tags on BBH tasks. Results show that the accuracy drops by 3 points when using 20 instead of 50 tags, whereas increasing the number of tags from 50 to 100 yields minimal change (see the Appendix for detailed results). This suggests that learning a sufficient – even redundant – number of tags can be beneficial for effectively expanding the output space.
@@ 6 Conclusion
We introduced fLSA, a foundation-model-based Latent Semantic Analysis method that aims to uncover the latent semantic structures in document collections by iteratively clustering and tagging document segments based on document-level contexts. Our experiments on story writing, math and multi-step reasoning tasks show that fLSA tags are more informative in reconstructing the original texts than tags generated by existing tagging methods. fLSA tags are also useful in expanding the output space via hierarchical sampling to increase the likelihood of discovering correct solutions to complex reasoning problems. These results suggest the potential of fLSA for generating effective task guidelines given some worked-out examples, along with hierarchical sampling and searching for problem solutions on challenging reasoning tasks.
7 Limitations
One limitation of fLSA is that some of the tags produced by fLSA may be semantically similar to each other, which can be ideally merged into a single tag. This limitation could be addressed by incorporating a tag fusion step in the EM algorithm, which we leave for future work. In addition, although the fLSA algorithm is agnostic to the LLM being used, we only test it on GPT-4 (which is one of the most powerful and widely used LLMs). Testing the algorithm on smaller models can be an interesting future work.
This work also has potential risks. One major risk is that the tags learned using fLSA may reflect the undesirable biases within the LLM being used. Integrating bias detection and mitigation techniques within the algorithm could be useful for addressing the issue.
Appendix A Appendix
A.1 Datasets
We evaluate fLSA against various baselines on story writing, math problem solving and multi-step reasoning benchmarks. We use WritingPrompts (Fan et al., 2018), a story writing dataset that contains 300K human-written stories paired with writing prompts from an online forum. We randomly sample 100 stories from the training set for clustering and tagging. We set the number of tags to 100 for all tagging approaches. For math problem solving, we use MATH (Hendrycks et al., 2021), a popular math benchmark that contains high school math competition problems on seven subjects including Prealgebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra and Precalculus. We learn 100 tags on 1K randomly sampled problem solutions from the training set. We also experiment on the Big-Bench Hard (BBH) benchmark (Suzgun et al., 2022). The original benchmark includes 23 challenging multi-step reasoning tasks, but each task only includes three step-by-step solution examples. Instead, we take the 12 tasks used in Xu et al. (2024) and learn the tags on the problem solutions (produced by their automatic prompt inference algorithm) for the 179 training problems. We set the number of tags to 50 for BBH. [^6] All datasets used in the work are under MIT license. Our use of the datasets is consistent with their intended use.
A.2 Large Language Model Setup
For clustering and tagging, we use GPT-4 (OpenAI et al., 2024) and Qwen-2.5-7B (a much smaller LLM introduced in Qwen et al. (2025)). For GPT-4, we set top_p=0.5top\_p=0.5, sampling temperature $\tau=1.0$, zero frequency and presence penalty. For Qwen-2.5-7B, we set top_p=0.5top\_p=0.5, sampling temperature $\tau=0.1$, zero frequency and presence penalty.
We also use GPT-4 with top_p=0.5top\_p=0.5 to estimate the reconstruction log-likelihood. We set the temperature $\tau=1.0$ when sampling alternative segments and $\tau=0$ when choosing the best continuation.
To measure Hits@K Accuracy, we use ChatGPT (gpt-3.5-turbo; OpenAI (2023)) instead of GPT-4. We set top_p=0.5top\_p=0.5 and temperature $\tau=1.0$ when sampling solutions from ChatGPT.
A.3 Ablation Study
5 shows the ablation study results on the number of tags.
20 Tags 50 Tags 100 Tags
Date 98.0 98.8 99.2
Formal 63.2 93.2 80.8
Geometric 86.4 87.6 86.0
Logical 98.9 99.5 99.1
Movie 93.6 95.2 94.8
ObjCount 99.6 95.2 99.6
Penguins 99.3 99.3 99.3
ReasonColored 100.0 100.0 100.0
RuinNames 90.8 93.6 95.6
TranslationError 72.8 75.2 72.4
Temporal 98.8 100.0 99.2
WordSort 57.6 56.0 61.6
Average 88.3 91.1 90.6
Table 5: Ablation Study: Hits@K Accuracy on BBH tasks using varying number of fLSA tags.